OpenStudy (anonymous):

whats (rad)3^2x+4=9^x-2

OpenStudy (anonymous):

rad?

OpenStudy (anonymous):

\[\huge \color{green}{\textbf{Welcome To Openstudy..}}\]

OpenStudy (mathteacher1729):

Can you type your equation in using the equation button?

OpenStudy (anonymous):

\[\sqrt{3}^{2x+4}=9^{x-2}\]

OpenStudy (anonymous):

9 = √3 ^ 4

OpenStudy (anonymous):

i know for that last question i have to set the exponents equal to each other

OpenStudy (anonymous):

yeaa

OpenStudy (anonymous):

but the base has to be same

OpenStudy (anonymous):

so whats x??

OpenStudy (anonymous):

ohhh mmmkkk i get ya thank!!!!!!!!!!!

OpenStudy (anonymous):

cool :)

OpenStudy (anonymous):

ill figure out x thanks for the help!!

OpenStudy (mathteacher1729):

The trick here is to "write nine as the square root of three ... raised to a power". Notice how I increase the power on each sqrt(3)... does this make sense, just from an exponents point of view? \[\huge \sqrt{3}^1=\sqrt{3}\] \[\huge \sqrt{3}^2=3\] \[\huge \sqrt{3}^3=3\cdot\sqrt{3}\] \[\huge \sqrt{3}^4=9\]

OpenStudy (anonymous):

thanks!!